OpenStudy (tia_sweetheart17):

solve the equation. x(5x+2)=3

OpenStudy (anonymous):

you will want to multiply the x through.. so x(5x+2) is 5x^2+2x = 3

OpenStudy (anonymous):

subtract 3 from both sides and you get 5x^2+2x-3=0

OpenStudy (anonymous):

do you know how to use the quadratic equation?

OpenStudy (anonymous):

x=-b +or- Sqrt(b^2-4ac) divided by 2a haha, does that make any sense @Tia_Sweetheart17

OpenStudy (anonymous):

you get that from ax^2+bx+c=0 so a=5 b=2 and c=-3

OpenStudy (anonymous):

I need to go, sorry and good luck!